# Inflationary power spectrum from the Lanczos algorithm

### v1

@@ V Bogoliubov transformation
Since the power spectrum of BD vacuum is well consistent with observations [6], we will use the Bogoliubov transformation to construct the relation between the BD vacuum and the vacuum of (15). First, the mode function $v_{z}(\eta,\vec{k})$ parametrizes the squeezed vacuum state in time-evolving spacetime. As for the reference BD vacuum mode, it is denoted by $v_{BD}(\eta,\vec{k})$. Then, we could introduce the non-trivial Bogoliubov transformation to relate $v_{z}(\eta,\vec{k})$ and $v_{BD}(\eta,\vec{k})$ as follows,
@@
Then, the we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
@@
where $\mathcal{N}$ represents the normalization factor, and $|0,0\rangle_{OST}$ denotes the vacuum of the OTMSS. The essence of Eq. (22) is that one can construct the wave function in the target spacetime using the Bogoliubov transformation and the BD vacuum. Therefore, the key aspect is determining the formulas for $\alpha_{k}$ and $\beta_{k}$ for our specific case. This is also crucial for understanding the relationship between the modes of the squeezed state and the modes of the BD vacuum, as indicated in Eq. (23). In fact, we can apply the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ to the annihilation operator of the BD vacuum to derive the coefficients for the Bogoliubov transformation,
@@
where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (15) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with [49, 50]. Next, we will proceed with the evolution of $\phi_{k}$ and $r_{k}$ determined by the Hamiltonian (11).
As we discussed, the OTMSS (15) is more generic compared with a pure two-mode squeezed state, including the dissipative effects of an open system. In this work, we will particularly focus on the super Hubble scale, namely, $k\eta$ will be approaching zero. In light of (13), we could see that $\mu_{2}$ (dissipative coefficient) will also be very tiny in this super Hubble scale, meaning that the universe within the super Hubble scale is a weak dissipative system. Consequently, the pure two-mode squeezed state will be a good approximation, and our derivation for (24) is also consistent with each other.
@@ VI Evolution of $\phi_{k}$ and $r_{k}$
To obtain the power spectrum of the OTMSS (15), one needs to rewrite the Bogoliubov coefficients (24) in terms of $k$ as in Refs. [51, 52, 53], where their formulas has made some approximation. To fully capture the information of (15), we will explicitly numerically give the evolution of $r_{k}$ and $\phi_{k}$ in terms of $k$, where the evolution is completely determined by the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. The resulting evolution equation of $r_{k}$ and $\phi_{k}$ can be denoted by via [26]
@@
Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential due to the slow-roll conditions. These calculations can be quite complicated; for more details, the reader can refer to the appendix of [26]. The evolution equations (26) for $r_{k}$ and $\phi_{k}$ describe the dynamics during the inflationary period. The power spectrum was observed at the horizon crossing at a fiducial momentum of $k_{*}=0.05~{}\text{Mpc}^{-1}$. To compare with observations, the data from the CMB covers the range from $10^{-4}~{}\text{Mpc}^{-1}$ to $1~{}\text{Mpc}^{-1}$ [6], making it beneficial to fully expand the power spectrum within this range. At the pivot point, we have $k_{*}=aH$, where the value is derived from today’s measurements, allowing us to easily obtain $\eta=-\frac{1}{k}$ within the CMB observational range [54, 55, 56]. Given that $\eta=-\frac{1}{aH}$, the evolution can be transformed accordingly,
$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(\mu_{2}^{\prime}+M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\\ &\ +M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (27)
Being armed with these two equations, we could give their numeric in terms of comoving momentum $k$.
@@
Fig. 3 clearly indicates that mean number density is almost zero since the particles are generated in preheating process, although the number density of particle is non-vanishing. Being armed with the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
@@ VII Power spectrum of OTMSS
Since the power spectrum of OTMSS (15) is proportional to $|v_{z}|^{2}$. Then, we can obtain the ratio between the power spectrum of the OTMSS and the BD vacuum in light of (23),
@@
where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of BD vacuum and we have used $\lim_{k\eta\to 0}\frac{v_{BD}^{*}}{v_{BD}}=-1$ (super Hubble scale approximation) via BD vacuum (5). Once having this relation, it is quite straightforwardly showing the numeric of (29) in terms of k.
@@
Fig. 4 indicates that the ratio (29) is almost one after the horizon crossing, inferring that the power spectrum of two-mode squeezed state will be consistent with the BD vacuum. For completeness, we will use the definition of power to confirm this conclusion. Following the notation of [34], the two-point correlation function can be defined as
@@
To match the observations, we will use the curvature perturbation to calculate the power spectrum, where the curvature perturbation is $\mathcal{R}=v/z$ with $z=\sqrt{2\epsilon}a$ as shown in Sec. II. Thus, we could obtain the power spectrum of curvature perturbation as follows,
@@
As for the power spectrum of BD vacuum, it will predict the $\Delta_{\mathcal{R}}^{2}=A_{s}(\frac{k}{k_{*}})^{n_{s}-1}$ where the observational constraints to the amplitude power spectrum $A_{s}$ and spectral index from [6]
@@
Figure 5: The numeric of squeezed parameter $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 5 clearly shows the power spectrum of the OTMSS is almost the same with BD vacuum due to the ratio relation (29) as shown in Fig. 4, meanwhile our evaluation for the spectral index $n_{s}$ is also consistent with the BD vacuum.
Recalling the calculation of the power spectrum for the squeezed state as outlined in [51, 52, 53], we can parametrize the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, in terms of $r_{k}$ and $\phi_{k}$. These coefficients will be expressed as functions of the momentum $k$. However, these procedures overlook a crucial aspect: both $r_{k}$ and $\phi_{k}$ evolve over time as determined by the corresponding Schrödinger equation, given by $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. Therefore, a more accurate approach would be to take the time evolution of $r_{k}$ and $\phi_{k}$ into consideration. Furthermore, we focus on the OTMSS, as described in (15), which represents the open system (the universe is regarded as an open system). Following this approach, our numerical results for the power spectrum are consistent with the BD vacuum. While our derivation cannot be distinguished from the BD vacuum, the wave function in (15) is more general, being model-independent and relying solely on the second Meixner polynomial and the Lanczos algorithm.
@@ VIII Conclusion and Discussion
Under the single field inflation, we have constructed a framework for calculating the power spectrum of two mode squeezed state, where the key ingredient is the OTMSS (15) and the generalized Lanczos algorithm as shown in Sec. III.
In light of the Lanczos algorithm, the Hamiltonian can be separated into components corresponding to open and closed systems, as shown in Eq. (14). During inflation, an amplification process drives quantum phenomena toward classicalization, making the two-mode squeezed state a natural framework for studying curvature perturbations. The squeezed state is constructed via the action of the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ on the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our case. Crucially, (14) reveals that $\mathcal{\hat{S}}_{-k,k}$ exclusively encodes the closed system dynamics. In contrast, the OTMSS in (15) incorporates environmental interactions through the dissipative coefficient $\mu_{2}$. The standard squeezed state emerges as the leading-order term in the weak dissipation limit ($\mu_{2}\to 0$). From the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$, we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, whose numerical solutions are plotted in Figs. 2 and 1. The ratio between the OTMSS and BD vacuum, defined via Bogoliubov transformations in Eq. (29). As shown in Fig. 4, $\gamma_{k}$ remains unity throughout inflation, consistent with suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics. Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in (15) matches the BD vacuum prediction, as demonstrated in Fig. 5.
@@
The derivation of the BD vacuum in (5) relies on solving the mode equation. In contrast, the OTMSS is constructed through the Hamiltonian formalism and Lanczos algorithm without solving equations of motion, where model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$. Remarkably, the power spectrum prediction from the OTMSS in (15) matches the BD vacuum result, suggesting that squeezed state construction is a more fundamental approach than solving EOM.
The key ingredient in this construction is the generalized displacement operator [57], defined as
@@
where $\hat{H}$ exhibits specific group structures (e.g., the Weyl algebra in our case) and the squeezed operator belongs this kind of operator. When $\hat{H}$ belongs to a Lie group, the operator $\mathcal{D}(\xi)$ provides a rigorous framework for generating quantum states, surpassing traditional EOM-based methods in universality.
@@
  • Current results apply to single-field inflation. Generalizations to multifield scenarios [58, 59, 60, 61, 62] and $f(R)$ gravity [63, 64] are natural next steps.
  • The $\mu_{2}$-dependence in OTMSS may encode decoherence effects observable in higher-order correlations.

### v2

@@ V Bogoliubov transformation
Since the power spectrum of the BD vacuum is in excellent agreement with observational data Planck:2018jri , we employ the Bogoliubov transformation to establish a relation between the BD vacuum and the vacuum state defined in Eq. (24). Specifically, the mode function $v_{z}(\eta,\vec{k})$ parameterizes the squeezed vacuum state in a time-dependent cosmological background. The reference mode function associated with the BD vacuum is denoted by $v_{\text{BD}}(\eta,\vec{k})$. We then introduce a non-trivial Bogoliubov transformation relating $v_{z}(\eta,\vec{k})$ and $v_{\text{BD}}(\eta,\vec{k})$ as follows:
@@
Then, we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
@@
where $\mathcal{N}$ is the normalization factor, and $|0,0\rangle_{\text{OST}}$ denotes the vacuum of the OTMSS, which is explicitly constructed from the orthogonal basis $|n,n\rangle_{-\vec{k},\vec{k}}$ generated via specific recurrence relations Gamayun:2025hvu ; Muck:2022xfc . The essence of Eq. (34) is that the wave function in the target spacetime can be systematically constructed using the Bogoliubov transformation applied to the BD vacuum. It is therefore essential to determine the explicit form of the Bogoliubov coefficients $\alpha_{k}$ and $\beta_{k}$ in our framework. This is crucial for understanding the correspondence between the squeezed state modes and those of the BD vacuum, as articulated in Eq. (35). In practice, we apply the two-mode squeezing operator
$\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$
to the annihilation operators of the BD vacuum to derive the Bogoliubov coefficients.
@@
where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (24) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with Bianchi:2024jmn ; Grain:2019vnq . Next, we will proceed with the evolution of $\phi_{k}$ and $r_{k}$ determined by the Hamiltonian (19).
As previously established, the OTMSS in Eq. (24) represents a generalization of the pure two-mode squeezed state, as it incorporates dissipative effects inherent to open quantum systems. In this work, we focus specifically on the super-Hubble regime, where $k\eta$ approaches zero. According to (22), $\mu_{2}$ (the dissipative coefficient) becomes extremely small in this regime, indicating that the universe at the super-Hubble scale behaves as a weakly dissipative system. Therefore, the pure two-mode squeezed state serves as a good approximation, and our derivation for (36) remains consistent with this understanding. In the sub-horizon regime, as $k\eta$ approaches infinity, the dissipative coefficient $\mu_{2}$ becomes the unity as discovered by Zhai:2024odw ; Li:2024kfm ; Li:2024iji ; Li:2024ljz This indicates that the sub-horizon region acts as a strongly dissipative system, leading to that we should use the OTMSS to investigate the observable by taking place of the two-mode squeezed state. Consequently, we expect observable predictions to deviate from those of the standard scenario. This provides a physical interpretation of observational consequences throughout the entire inflationary epoch: initially, observable outcomes deviate from the standard case in the sub-horizon region. After horizon exit, when the mode enters the super-horizon regime, the observational signatures align with the standard results. In the following calculations, we will use the power spectrum to illustrate this point in the super-horizon regime since it is related to the observations.
@@ VI Evolution of $\phi_{k}$ and $r_{k}$
To derive the power spectrum of the OTMSS given in Eq. (24), the Bogoliubov coefficients in Eq. (36) must be expressed as functions of the wavenumber $k$, following the approach adopted in Refs. Danielsson:2002kx ; Danielsson:2004xw ; Broy:2016zik , where certain approximations were employed. To fully incorporate the structure of the OTMSS wave function (24), we numerically compute the evolution of the squeezing parameters $r_{k}$ and $\phi_{k}$ with respect to $k$. This evolution is governed by the Schrödinger equation
$\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle,$
which completely determines the dynamics. The resulting equations of motion for $r_{k}$ and $\phi_{k}$ can be expressed as derived in Zhai:2024odw :
@@
Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential term due to the slow-roll conditions. These calculations can be rather involved; for complete details, we refer the reader to the appendix of Zhai:2024odw . The evolution equations (38) for $r_{k}$ and $\phi_{k}$ describe the squeezing dynamics during the inflationary epoch. The power spectrum is evaluated at horizon crossing for a fiducial scale $k_{*}=0.05~\text{Mpc}^{-1}$. To facilitate comparison with observations, we note that CMB measurements cover the range from $10^{-4}~\text{Mpc}^{-1}$ to $1~\text{Mpc}^{-1}$ Planck:2018jri . It is therefore useful to compute the power spectrum across this full range. Under these considerations, Eq. (38) becomes:
$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (39)
To obtain the corresponding numerical values of $r_{k}$ and $\phi_{k}$, we use the horizon-crossing condition $k_{*}=aH$ at the pivot scale, whose value is inferred from present-day observations. This allows us to express the conformal time as $\eta=-1/k$ across the CMB observational range Bianchi:2024qyp ; Gong:2001he ; Auclair:2022yxs . Using these evolution equations, we numerically compute $r_{k}$ and $\phi_{k}$ as functions of the comoving momentum $k$.
@@
Fig. 3 clearly indicates that mean number density is almost zero since the particles are generated in preheating process, although the number density of particle is non-vanishing. Using the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
@@ VII Power spectrum of OTMSS
Since the power spectrum of the OTMSS in Eq. (24) scales with $|v_{z}|^{2}$, we may express the ratio between this power spectrum and that of the BD vacuum as follows, using the Bogoliubov transformation given in Eq. (35):
@@
where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of the BD vacuum, and we have used the super-Hubble scale approximation $\lim_{k\eta\to 0}v_{BD}^{*}/v_{BD}=-1$ derived from the BD vacuum definition in Eq. (5). With this relation established, it is straightforward to compute the numerical evaluation of Eq. (41) as a function of $k$.
@@
Fig. 4 shows that the ratio in Eq. (41) is very close to unity after horizon crossing, which implies that the power spectrum of the two-mode squeezed state is consistent with that of the BD vacuum. For completeness, we further validate this result using the formal definition of the power spectrum. Following the notation of Baumann:2009ds , the two-point correlation function is defined as
@@
To compare with observational data, we compute the power spectrum using the curvature perturbation $\mathcal{R}=v/z$, where $z=\sqrt{2\epsilon}\,a$ as defined in Sec. II. The power spectrum of the curvature perturbation is therefore given by:
@@
As for the power spectrum of the BD vacuum, it takes the form
$\Delta_{\mathcal{R}}^{2}=A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},$
where the amplitude $A_{s}$ and the spectral index $n_{s}$ are constrained by observational data from Planck:2018jri .
@@
Figure 5: The numeric of power spectrum $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 5 clearly demonstrates that the power spectrum of the OTMSS is nearly identical to that of the BD vacuum, which is consistent with the near-unity ratio shown in Eq. (41) and Fig. 4. Furthermore, our evaluation of the spectral index $n_{s}$ agrees with the BD vacuum results within observational uncertainties.
Recalling the calculation of the power spectrum for the squeezed state as outlined in Danielsson:2002kx ; Danielsson:2004xw ; Broy:2016zik , we may parametrize the Bogoliubov coefficients $\alpha_{k}$ and $\beta_{k}$ in terms of the squeezing parameters $r_{k}$ and $\phi_{k}$, which are in turn expressed as functions of the momentum $k$. However, these treatments neglect an essential feature: both $r_{k}$ and $\phi_{k}$ evolve in time according to the Schrödinger equation. A more complete approach must therefore incorporate this dynamical evolution. Moreover, we focus on the OTMSS defined in Eq. (24), which inherently describes an open quantum system—reflecting the fact that the universe is best treated as an open system. Within this framework, our numerical results for the power spectrum align with those of the BD vacuum. Although our predictions are observationally indistinguishable from the BD vacuum, the wave function in Eq. (24) is more general: it is model-independent and constructed solely from the second kind of Meixner polynomial within the generalized Lanczos algorithm.
@@ VIII Conclusion and Discussion
Within the framework of single-field inflation, we have developed a method for computing the power spectrum of the two-mode squeezed state. The central elements of this construction are the OTMSS given in Eq. (24) and the generalized Lanczos algorithm, as detailed in Sec. III.
Within the framework of the generalized Lanczos algorithm, the Hamiltonian can be decomposed into open and closed system components, as expressed in Eq. (23). During inflation, an amplification mechanism drives the quantum-to-classical transition, making the two-mode squeezed state a natural basis for characterizing curvature perturbations. This squeezed state is constructed by applying the squeezing operator
$\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$
to the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our notation. Crucially, Eq. (23) indicates that $\mathcal{\hat{S}}_{-k,k}$ captures only the closed system dynamics. In contrast, the OTMSS defined in Eq. (24) incorporates environmental interactions via the dissipative coefficient $\mu_{2}$. The conventional squeezed state is recovered in the weak dissipation limit ($\mu_{2}\to 0$) as the leading-order approximation.
From the Schrödinger equation
$\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle,$
we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, with numerical solutions illustrated in Figs. 2 and 1. The ratio $\gamma_{k}$ between the OTMSS and BD vacuum power spectra, defined via Bogoliubov transformations in Eq. (41), remains unity throughout inflation, as shown in Fig. 4. This is consistent with strongly suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics. Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in Eq. (24) matches the prediction of the BD vacuum, as confirmed in Fig. 5.
@@
The derivation of the BD vacuum in Eq. (5) relies on explicitly solving the mode equation of motion. In contrast, the OTMSS is systematically constructed within the Hamiltonian formalism via the Lanczos algorithm without solving the equations of motion; instead, model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$. Remarkably, the power spectrum predicted by the OTMSS in Eq. (24) matches the result from the BD vacuum, indicating that the squeezed-state formulation provides a more fundamental description than direct calculation of the equations of motion.
The key ingredient in this construction is the generalized displacement operator Caputa:2021sib , defined as
@@
where $\hat{H}$ exhibits specific group structures (e.g., the $SL(2.R)$ in our case) and the squeezed operator belongs this kind of operator. When $\hat{H}$ belongs to a Lie group, the operator $\mathcal{D}(\xi)$ provides a rigorous framework for generating quantum states, surpassing traditional EOM-based methods in universality.
@@
  • Current results apply to single-field inflation. Generalizations to multifield scenarios Liu:2019xhn ; Liu:2020zzv ; Liu:2020zlr ; Liu:2021rgq ; Zhang:2022bde and $f(R)$ gravity Liu:2018hno ; Liu:2018htf are natural next steps.
  • The $\mu_{2}$-dependence in OTMSS may encode decoherence effects observable in higher-order correlations. Pronounced dissipative effects on sub-horizon scales are expected to cause deviations in non-Gaussian correlators from those predicted by the standard two-mode squeezed state. These significant dissipative effects may also influence decoherence.